A Bibliography of Publications of Alexander Strehl

Alexander Strehl The University of Texas at Austin

Austin, TX, USA

Tel: ?n/a? FAX: ?n/a?

E-mail: alexander@strehl.com (Internet)

09 April 2018 Version 1.02

Abstract

This bibliography records publications of Alexander Strehl.

Title word cross-reference

Affine [Str98]. Airborne [SA99, SA00a]. Algorithm [Str98]. Approach [SG00b]. Association [GSG99].

Balanced [SG00b]. Based [GSG99, SG03, SG00a, SA00a, SG00c, SG01]. baskets [SG00b]. Bayesian [Str98, SA00b].

Clustering [GSG99, SG00a, SGM00, SG00b, SG03]. Clusters [SG01]. Customer [SG00c].

Data [SG00c, SG01, SG03]. Data-sets [SG00c]. Detecting [SA99]. Detection [SA00a]. Dimensional [SG03, SG00b, SG01]. Distance [GSG99].

Estimation [Str98, SA00a, SA00b]. Evaluation [SG00a].

FLIR [SA00a]. Forward [SA99]. Framework [SA00b].

Grouping [SG00c]. Guidance [SG00a].

High [SG00b, SG01, SG03]. High-Dimensional [SG03, SG00b, SG01].

Impact [SGM00]. Infra [SA99]. Infra-Red [SA99].

Large [SG00c]. Looking [SA99].

Market [SG00b]. Market-baskets [SG00b]. Measures [SGM00]. Method [SA00a]. Mining [SG03]. MODEEP [SA00a]. Motion [Str98, SA00a]. Motion-based [SA00a]. Motions [Str98, SA00b]. Moving [SA99]. Multiple [SA00b].

Object [SA00a]. Objects [SA99]. page [SGM00]. Pose [SA00a]. Presence [Str98].

Quality [SG00a].

Red [SA99]. Relationship [SG00a, SG01, SG03]. Relationship-Based [SG03, SG00a, SG01]. Relaxation [Str98, SA00b]. Retail [SG00c]. Rules [GSG99].

Scalable [SG00b]. Segmentation [Str98, SA00b]. Sequences [SA99, SA00a]. sets [SG00c]. Similarity [SGM00].

Two [Str98].

Using [SG00a].

Value [SG00c]. Value-based [SG00c]. Visualization [SG00a, SG01, SG03].

Web [SGM00]. Web-page [SGM00].

References

Gupta:1999:DBC

[GSG99] Gunjan K. Gupta, Alexander Strehl, and Joydeep Ghosh. Distance based clustering of association rules. In Proceedings of ANNIE 1999 (Intelligent Engineering Systems Through Artificial Neural Networks), 7–10 November 1999, St. Louis, Missouri, USA, volume 9, pages 759– 764. ASME, November 1999.

Strehl:1999:DMO

[SA99] Alexander Strehl and J. K. Aggarwal. Detecting moving objects in airborne forward looking infra-red sequences. In Proceedings of the 1999 IEEE Workshop on Computer Vision Beyond the Visible Spectrum: Methods and Applications (CVBVS 1999), 21–22 June 1999, Fort Collins, Colorado, USA, pages 3–12. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 1999.

Strehl:2000:MMB

[SA00a] Alexander Strehl and J. K. Aggarwal. MODEEP: A motion-based object detection and pose estimation method for airborne FLIR sequences. Machine Vision and Applications, 11(6):267–276, 2000. CO-DEN MVAPEO. ISSN 0932-8092.

Strehl:2000:NBR

[SA00b] Alexander Strehl and J. K. Aggarwal. A new Bayesian relaxation framework for the estimation and segmentation of multiple motions. In Proceedings of the 4th IEEE Southwest Symposium on Image Analysis and Interpretation (SSIAI 2000), 2–4 April 2000, Austin, Texas, USA, pages 21–25. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 2000.

Strehl:2000:CGQ

[SG00a] Alexander Strehl and Joydeep Ghosh. Clustering guidance and quality evaluation using relationshipbased visualization. In Proceedings of ANNIE 2000 (Intelligent Engineering Systems Through Artificial Neural Networks), 5–8 November 2000, St. Louis, Missouri, USA, volume 10, pages 483–488. ASME, November 2000.

Strehl:2000:SAB

[SG00b] Alexander Strehl and Joydeep Ghosh. A scalable approach to balanced, high-dimensional clustering of market-baskets. In Proceedings of the Seventh International Conference on High Performance Computing (HiPC 2000), 17–20 December 2000, Bangalore, India, volume 1970 of Lecture Notes in Computer Science, pages 525–536. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., December 2000.

Strehl:2000:VBC

[SG00c] Alexander Strehl and Joydeep Ghosh. Value-based customer grouping from large retail data-sets. In Proceedings of the SPIE Conference on Data Mining and Knowledge Discovery: Theory, Tools, and Technology II, 24–25 April 2000, Orlando, Florida, USA, volume 4057, pages 33–42. SPIE, April 2000.

Strehl:2001:RBV

[SG01] Alexander Strehl and Joydeep Ghosh. Relationship-based visualization of high-dimensional data clusters. In Proceedings of KDD Workshop on Visual Data Mining, August 26, 2001, San Francisco, pages 90–99. ????, ????, August 2001.

Strehl:2003:RBC

[SG03] Alexander Strehl and Joydeep Ghosh. Relationship-based clustering and visualization for highdimensional data mining. IN-FORMS Journal on Computing, 15 (2):208–230, Spring 2003. CODEN ???? ISSN 1091-9856 (print), 1526-5528 (electronic). URL https:

                                        //pubsonline.informs.org/doi/
                                        abs/10.1287/ijoc.15.2.208.14448.

Strehl:2000:ISM

[SGM00] Alexander Strehl, Joydeep Ghosh, and Raymond Mooney. Impact of similarity measures on web-page clustering. In Proceedings of the 17th National Conference on Artificial Intelligence: Workshop of Artificial Intelligence for Web Search (AAAI 2000), 30–31 July 2000, Austin, Texas, USA, pages 58–64. AAAI, July 2000.

Strehl:1998:NBR

[Str98] Alexander Strehl. A new Bayesian relaxation algorithm for motion estimation and segmentation in the presence of two affine motions. Master’s thesis, The University of Texas at Austin, Austin, TX, USA, August 1998.